OpenStudy (anonymous):

Ten flight attendants based in New York and six flight attendants based in Boston have formed a flight regulations oversight committee. A subcommittee made up of five people from the group will delegate the work to be done. What are the odds that the subcommittee will contain three flight attendants from New York and two from Boston? A. 18:182 B. 225:308 C. 75: 107 D. 18:25

OpenStudy (anonymous):

I got D

OpenStudy (kropot72):

It is not D. Do you want help to work through it?

OpenStudy (kropot72):

The probability that 3 will be from NY is given by: \[P(3\ from\ NY)=\frac{\left(\begin{matrix}10 \\ 3\end{matrix}\right)\left(\begin{matrix}6 \\ 2\end{matrix}\right)}{\left(\begin{matrix}16 \\ 5\end{matrix}\right)}=\frac{75}{182}\] P(other than 3 from NY)=1 - (75/182) The required odds are the ratio of the probability of '3 from NY' to the probability of 'other than 3 from NY'. Do you need help to finish?

OpenStudy (anonymous):

yes i need help to finish it.

OpenStudy (anonymous):

It is D

OpenStudy (anonymous):



OpenStudy (anonymous):

but the person it wasnt d... are you sure its d?

OpenStudy (kropot72):

P(3 from NY) = 75/182 P(other than 3 from NY)=1 - (75/182) = 107/182. The required odds are therefore = 75/182 : 107/182 = 75 : 107 The correct choice is C.

OpenStudy (anonymous):

i just recalculated it and it was c